SIMULATION OF THE 22μ INFRARED HOT BANDS OF NITRIC ACID FROM THE ANALYSES OF ROTATIONAL SPECTRA

\underline{DOUGLAS T. PETKIE}, Department of Physics, Wright State University, Dayton OH 45435; PAUL HELMINGER, Department of Physics, University of South Alabama, Mobile, AL 36688; BRENDA P. WINNEWISSER, MANFRED WINNEWISSER, and FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210.

Resent results of millimeter and submillimeter wave rotational spectroscopy of the ν_{5}/2ν_{9} states of nitric acid^{a} are used to simulate the complex structure of the 2ν_{9} − ν_{9} and ν_{5} − ν_{9} hot bands near 22μ. The comparison data were obtained with a high resolution Bruker FTIR at various pressures and a temperature of 297 K. The combination of the quality and resolution of the infrared data and the complexity of the spectra of these interacting states represents a stringent test for the simulation. It is shown that the agreement is very good and that this approach is generally advantageous.

aD. T. Petkie, T. M. Goyette, P. Helminger, H. M. Pickett, and F. C. De Lucia, J. Mol. Spectrosc. 208, 121-135 (2001).